# Nearest Neighbor Analysis as a New Probe for Fuzzy Dark Matter

### v1

@@ Acknowledgements
We would like to express our appreciation to Shant Baghram for the insightful discussions that greatly influenced this work. We are also grateful to Simon May and Volker Springel for providing us with the halo catalogs of their large-volume simulations. We acknowledge using yt Project, GriSPy package and SciPy library for data analysis and making plots. Finally, we would like to acknowledge that MA is being supported by a grant from the Iran National Elite Foundation (INEF) under the Chamran Plan for Young Researchers.
Appendix A Supplementary Figures
@@
uqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@ Acknowledgements
We would like to express our appreciation to Shant Baghram for the insightful discussions that greatly influenced this work. We are also grateful to Simon May and Volker Springel for providing us with the halo catalogs of their large-volume simulations. We acknowledge using yt Project, GriSPy package, and SciPy library for data analysis and making plots. Finally, we would like to acknowledge that MA is being supported by a grant from the Iran National Elite Foundation (INEF) under the Chamran Plan for Young Researchers.
Data Availability
The halo catalogs for the TNG simulations can be accessed at https://www.tng-project.org/data/. On the other hand, the halo catalogs for the simulations performed by May & Springel (2021) are available upon request to them. Furthermore, we will be happy to share the results of our analysis upon a reasonable request.
Appendix A Redshift Dependence
Table 4: Moments of SC-PDF at different redshifts
Simulation redshift $s_{1}$ $s_{2}$ $s_{3}$ $s_{4}$
CDM 3 $1.563\pm 0.016$ $0.876\pm 0.015$ $0.925\pm 0.043$ $3.92\pm 0.168$
CDM 5 $1.517\pm 0.015$ $0.848\pm 0.013$ $0.931\pm 0.036$ $3.942\pm 0.114$
CDM 7 $1.553\pm 0.015$ $0.859\pm 0.013$ $0.883\pm 0.042$ $3.814\pm 0.143$
FDM (CDM IC), $8640^{3}$, $m_{22}=0.7$ 3 $1.134\pm 0.013$ $0.547\pm 0.007$ $0.789\pm 0.024$ $3.577\pm 0.1$
FDM (CDM IC), $8640^{3}$, $m_{22}=0.7$ 5 $1.064\pm 0.014$ $0.492\pm 0.007$ $0.689\pm 0.012$ $3.37\pm 0.043$
FDM (CDM IC), $8640^{3}$, $m_{22}=0.7$ 7 $1.061\pm 0.015$ $0.484\pm 0.008$ $0.687\pm 0.03$ $3.32\pm 0.075$
Table 5: Logarithmic moments of NN-PDF at different redshifts
Simulation redshift $l_{1}$ $l_{2}$ $l_{3}$ $l_{4}$
CDM 3 $-0.988\pm 0.008$ $0.6\pm 0.003$ $0.363\pm 0.006$ $3.038\pm 0.012$
CDM 5 $-0.964\pm 0.006$ $0.58\pm 0.003$ $0.492\pm 0.006$ $3.097\pm 0.015$
CDM 7 $-1.039\pm 0.005$ $0.601\pm 0.004$ $0.544\pm 0.006$ $3.196\pm 0.024$
FDM (CDM IC), $8640^{3}$, $m_{22}=0.7$ 3 $-0.555\pm 0.004$ $0.503\pm 0.004$ $-0.144\pm 0.024$ $3.113\pm 0.037$
FDM (CDM IC), $8640^{3}$, $m_{22}=0.7$ 5 $-0.457\pm 0.007$ $0.468\pm 0.004$ $-0.275\pm 0.026$ $3.046\pm 0.056$
FDM (CDM IC), $8640^{3}$, $m_{22}=0.7$ 7 $-0.422\pm 0.01$ $0.448\pm 0.005$ $-0.221\pm 0.043$ $3.207\pm 0.086$
Table 6: Nearest Neighbors’ Angle Analysis Moments at different redshifts
Simulation redshift Mean of $\phi$ Std of $\phi$ $a_{4}$
CDM 3 $81.51\pm 0.13^{\circ}$ $46.49\pm 0.22^{\circ}$ $0.9623\pm 0.0375$
CDM 5 $82.86\pm 0.11^{\circ}$ $48.18\pm 0.17^{\circ}$ $1.1168\pm 0.0252$
CDM 7 $82.31\pm 0.22^{\circ}$ $49.53\pm 0.09^{\circ}$ $1.3591\pm 0.0219$
FDM (CDM IC), $8640^{3}$, $m_{22}=0.7$ 3 $84.93\pm 0.28^{\circ}$ $43.93\pm 0.17^{\circ}$ $0.6234\pm 0.0908$
FDM (CDM IC), $8640^{3}$, $m_{22}=0.7$ 5 $85.43\pm 0.38^{\circ}$ $42.79\pm 0.28^{\circ}$ $0.4421\pm 0.0507$
FDM (CDM IC), $8640^{3}$, $m_{22}=0.7$ 7 $84.27\pm 0.41^{\circ}$ $40.97\pm 0.2^{\circ}$ $-0.2033\pm 0.0513$
In the main text, we compared the simulations only at redshift $z=3$. In this appendix, we investigate whether our results vary at higher redshifts. Specifically, we repeat our analysis for the "CDM" and "FDM (CDM IC), $8640^{3}$, $m_{22}=0.7$" simulations at redshifts $z=5$ and $z=7$, and compare them with the results obtained at $z=3$. In Figs. 8(a)-(b) we observe that the distribution of SC and NND in the FDM also deviates from the CDM at higher redshifts $z=5$ and $z=7$, which is consistent with our previous results at $z=3$. The deviation of the FDM from the CDM at higher redshifts is also observable in the PDFs of the NNA, which are shown in Figs. 8(c)-(d). Interestingly, the deviation becomes larger as we look at the higher redshifts. This can be easily seen by looking at the lowest and highest $\phi$’s and $\cos\phi$’s in Figs. 8(c)-(d).
Tables. 4-6, list the moments of the SC-PDF, the logarithmic moments of the NN-PDF, and the moments of the NNA analysis for CDM and FDM simulations at different redshifts. In almost all of them, we see a monotonic increase in the deviation between CDM and FDM as we go to higher redshifts. This is consistent with some previous studies suggesting that the statistical properties of the FDM deviate less from the CDM at lower redshifts(see, e.g., Nori et al., 2018; higherredshifts; Ferreira, 2021; HMFofFDM). This is partly due to the coupling between the long and short modes during the nonlinear evolution, and also to the shrinking of the FDM cutoff scale at lower redshifts.
(a) PDF of SC
(b) PDF of NND
(c) PDF of NNA ($P(\phi)$)
(d) PDF of NNA ($P(\cos\phi)$)
Figure 8: The SC-PDF, NND-PDF, and NNA-PDF as a function of $\cos\phi$ and $\phi$ at redshifts $z=3$, $z=5$, and $z=7$ for the CDM simulation and the FDM simulation with CDM IC and $m_{22}=0.7$. The deviation of FDM from CDM increases at higher redshifts.
Appendix B Supplementary Figures